Circular No. 7598

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001af IN MCG –04-24-1 M. Modjaz and W. D. Li, University of California at Berkeley, on behalf

of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 16.8) on unfiltered CCD images taken on Mar. 18.3 and 19.3 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001af is located at α = 9^{h}50^{m}28^{s}.70, δ = −21^{o}48′30′′.7 (equinox 2000.0), which is 13′′.6 east and 22′′.7 south of the nucleus of MCG –04-24-1. An unfiltered KAIT image of the same field on Mar. 17.3 already showed a hint of the new object, while an image on Mar. 1.3 showed nothing at this position (limiting mag ∼ 18.5).

1998 SF_{36}

R. P. Binzel and A. S. Rivkin, Massachusetts Institute of Technology, report that they obtained spectra of 1998 SF_{36} (cf. MPEC 1998-S45, MPEC 2000-U40, MPEC 2001-F01, MPO 10868), a candidate target for the Japanese MUSES-C sample-return mission, at 0.5–1.0 μm. Measurements, obtained on Mar. 6.4 UT using the Kitt Peak 4-m reflector, reveal an S-type asteroid reflectance.

M. Hicks, P. Weissman, A. Chamberlin, and S. Lowry, Jet Propulsion Laboratory, write: “We have obtained low-resolution CCD spectroscopy of 1998 SF_{36} from 0.35 to 1.0 μm, using the Palomar 5-m Hale reflector on Mar. 17.43 UT. The object exhibits a spectrum dominated by olivine. The modest spectral slope between 0.55 and 0.70 μm (S = 8.0 ± 0.3 percent per 100 nm) and deep 1-μm absorption (V − X = +0.19 ± 0.003) suggests a classification of type QRS, similar to other small basaltic near-earth objects in the ECAS dataset (Cruikshank et al. 1991, Icarus 89, 1).”

T. Sekiguchi, M. Sterzik, N. Ageorges, and O. Hainaut, European Southern Observatory (ESO), communicate: “Our thermal observations (total integration time 193.6 s, using the ESO 3.6-m telescope + TIMMI2 at La Silla) of 1998 SF_{36} on Mar. 14.24 UT yield a preliminary measurement of the N-band (11.9 μm) radiation of 0.26 ± 0.03 Jy. Assuming the standard thermal model with beaming parameter 1.2, slope parameter 0.15 for near-earth objects, and absolute (V) magnitude of H = 19.1, the derived diameter is 0.36 ± 0.02 km and the geometric albedo is 0.32 ± 0.04. While these values are model-dependent, they clearly show that 1998 SF_{36} is not large.”

2001 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green